OpenStudy (anonymous):

Show me how to walk through (4x=3) + (5x-2)=55

OpenStudy (anonymous):

(4x+3) + (5x-2) = 55 9x +1 = 55 9x = 54 x=54/9 x = 6

OpenStudy (anonymous):

assuming ur '=' is to be a '+' (4x+3) + (5x-2)=55 9x +1=55 (by collecting like terms) 9x=54 x=9